# Standard modules of affine Hecke algebras

### v1

@@ 3.2 Proof of Theorem 2 for $G=\mathrm{GL}_{n}$
    Here, the first equality used that the intersection pairing commutes with pullback by the vertical flat maps in the diagram
    $\gamma(\xi)$$\tilde{\mathcal{N}}^{\gamma(s)}\times\tilde{\mathcal{N}}^{\gamma(s)}$$\xi$$\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}}}$
    by [Fulton, Thm. 6.2 (b)]. The second equality follows from commutativity of the obvious square, viewing $v$ as a cycle in the ambient smooth variety $\tilde{\mathcal{N}}^{s}$.
The proof of Theorem 2 for $G=\mathrm{GL}_{n}$ is now complete. ∎
@@
In general, the action of $\mathcal{L}_{\sigma}$ on $H_{*}(\mathcal{B}_{N}^{s};\mathbb{C})$ depends only on the conjugacy class of Cartan subgroup of (the Levi quotient of) the disconnected group $Z_{G}(N)$, see [CellsIV, 3.4 (a)], or [rigid, Lem. 14, §3.5] for this observation in the language of the operator Paley-Wiener theorem for Hecke algebras. In either perspective, either because $Z_{\mathrm{GL}_{n}}(N)$ is connected, or because the only $I$-spherical discrete series of each Levi subgroup of $\mathrm{GL}_{n}$ is the tensor product of Steinberg representations, respectively, it follows that the action depends only on $N$ in this case.

### v2

@@ 3.2 Proof of Theorems 1 and 2 for inner forms of $\mathrm{GL}_{n}$
    Here, the first equality used that the intersection pairing commutes, by [Fulton, Thm. 6.2 (b)], with pullback by the vertical flat maps in the diagram
    $\gamma(\xi)$$\tilde{\mathcal{N}}^{\gamma(s)}\times\tilde{\mathcal{N}}^{\gamma(s)}$$\xi$$\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}.$$\scriptstyle{\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}}}$
    The second equality follows from commutativity of the obvious square, viewing $v$ as a cycle in the ambient smooth variety $\tilde{\mathcal{N}}^{s}$.
The proof of Theorems 1 and 2 for inner forms of $\mathrm{GL}_{n}$ is now complete. ∎
@@
In general, the action of $\mathcal{L}_{\sigma}$ on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{C}}$ depends only on the conjugacy class of Cartan subgroup of (the Levi quotient of) the disconnected group $Z_{G}(N)$; see [CellsIV, 3.4 (a)], or [rigid, Lem. 14, §3.5] for this observation in the language of the operator Paley-Wiener theorem for Hecke algebras. In either perspective, either because $Z_{\mathrm{GL}_{n}}(N)$ is connected, or because the only $I$-spherical discrete series of each Levi subgroup of $\mathrm{GL}_{n}$ is the tensor product of Steinberg representations, respectively, it follows that the action depends only on $N$ in this case.